OpenStudy (anonymous):

24. The perimeter of a rectangular concrete slab is 114 feet and its area is 702 square feet. Find the dimensions of the rectangle. a. Using l for the length of the rectangle, write an expression for the width of the rectangle in terms of l. (Hint: Solve the formula for w.) Show your work. b. Write a quadratic equation using l, the expression you found in part (a), and the area of the slab. c. Solve the quadratic equation. Use the two solutions to find the dimensions of the rectangle.

OpenStudy (radar):

a. 2l + 2w = 114 2w = 114-2l w=57 - l Any questions on the part a.??

OpenStudy (radar):

Part b. Following the instructions provided with the problem. A = lw A=l(57-l) A = 57l - l^2 702= 57l-l^2 -l^2 + 57l-702 = 0 l^2 -57l + 702 = 0 That completes part b. Since you have no questions, Leaving Part C. to you.